Predict the outcome of the following cross: Hh IAIO x Hh IBIO

* refer the Bombay phenotype
Answer #1

Answer:

H antigen deficiency is known as the "Bombay phenotype" (h/h, also known as Oh): If H antigen is not present, no conventional blood types are recognized.

Hh IAIO x Hh IBIO

HIB

HIO

hIB

hIO

HIA

HIA / HIB (AB-type)

HIA / HIO(A-type)

HIA/ hIB(AB-type)

HIA/ hIO(A-type)

HIO

HIO/ HIB(B-type)

HIO / HIO(O-type)

HIO/ hIB(B-type)

HIO/ hIO(O-type)

hIA

hIA/ HIB(AB-type)

hIA / HIO(A-type)

hIA/ hIB (Bombay-type)

hIA/ hIO(Bombay-type)

hIO

hIO/ HIB(B-type)

hIO /HIO(O-type)

hIO/ hIB(Bombay-type)

hIO/ hIO(Bombay-type)

AB-type = 3/16

A-tpe = 3/16

B-type = 3/16

O-type = 3/16

Bombay-type = 4/16